# Modeling fibrous tissue in vascular fluid-structure interaction: a morphology-based pipeline and biomechanical significance

### v1

@@ 3.3 Patient-specific study
The images at the cross-sections (illustrated by the dashed lines in Figure 18 (a)) of the visceral arteries, supra-celiac (SC) aorta, and infrarenal (IR) aorta were obtained using a pulse and respiratory gated 2D phase contrast sequence. The obtained 2D phase-contrast magnetic resonance imaging (PCMRI) data were processed using Segment [1] to extract the physiological flow rates at these cross-sections for 30 time frames within a cardiac cycle. Polynomial interpolation and phase correction were applied to these flow rates to obtain the flow waveforms at the inlet and outlets through a cardiac cycle. Due to limitations in spatial resolution, the 2D PCMRI technique is insufficient to accurately identify small arteries, such as the left gastric arteries and lumbar arteries, their flow rates were excluded from this study. To ensure mass conservation, the sum of the flow waveforms of the IR aorta and all visceral arteries was set as the flow waveform for the inlet of the SC aorta. These obtained flow waveforms were then combined with a parabolic flow profile to prescribe the velocity at the inlet and the outlets of the five visceral arteries (Figure 18 (b)).
@@
Figure 20: The mesh of AA model.
Regarding the outlet of the IR aorta, a three-element Windkessel model was coupled with the 3D FSI model as the outflow boundary condition. Utilizing the measured flow waveform at the IR arterial outlet, the RCR parameters were tuned to ensure that the peak and trough of the output pressure matched the systolic pressure ($P_{\mathrm{sys}}=103$ mmHg) and diastolic pressure ($P_{\mathrm{dia}}=72$ mmHg) of the subject. The specific settings are as follows. The compliance of the Windkessel model was set to 0.000452 cm3/dyn; the proximal and distal resistances were set to 330.69 $\cdots$/cm3 and 7708.28 $\cdots$/cm3, respectively. Additionally, in this study, the exterior wall of the artery was kept stress-free, meaning the supporting tissue was not considered. Meanwhile, on the inlet and outlet solid surfaces, the ring-shaped walls were fixed.
We employed the isotropic neo-Hookean model (2.4) and the anisotropic GOH model (2.5) for the arterial wall in the FSI simulations, respectively. The solid material properties utilized were adopted from [9] and are summarized in Table 1. Moreover, the effects of fiber dispersion was investigated by considering two different dispersion parameters (i.e., $\kappa=0.0$ and $0.226$ [38]). The blood density was set to 1.06 g/cm3, and the dynamic viscosity was taken to be 0.04 $\cdots$/cm2. The geometric model was discretized into $8.52\times 10^{6}$ linear tetrahedral elements, including $4.64\times 10^{6}$ solid and $3.89\times 10^{6}$ fluid elements. Three layers of BL elements with a growth ratio of 0.75 were constructed within the lumen. In the meantime, we strived to have at least six layers of elements across the vessel wall thickness (Figure 20). We also investigated the problem with a relatively coarser mesh, which delivered essentially identical results. The following reported results can be viewed as mesh independent. To construct the orientation of fibers, the local basis vectors were generated on the vessel wall using the procedures proposed in Section 2.2 (Figure 19).
@@
Figure 21: The velocity computed near the peak systole.
FSI simulations were performed using the initialization procedure outlined in Section 2.4. Two cardiac cycles were simulated with a fixed time step size of $\Delta t=1.0\times 10^{-5}$ s, and the results from the last cycle are presented here. Figures 21 and 22 depict the flow fields and WSS obtained by the four different material models at peak systole. It is evident that the flow and WSS distributions derived from these models are quite close, with only minor discrepancies in the flow field near the outlet of the IR aorta. Consequently, when focusing solely on the flow field and WSS, it is acceptable to approximate the arterial wall as an isotropic material.
Figure 22: The WSS computed near the peak systole.
Figure 23: The maximum principal strain of the arterial wall computed near the peak systole.
Figures 23 and 24 depict the maximum principal strain and von Mises stress at peak systole for four material models, respectively. Significant differences are observed between isotropic and anisotropic models. For the isotropic models, the principal strain is lower and distributed smoothly within the wall; for the anisotropic models, the principal strain is higher in the visceral arteries. The stress pattern exhibits similarity among the two neo-Hookean and the GOH model with $\kappa=0.226$. Comparing the stress results for the GOH model with $\kappa=0.0$ and $\kappa=0.226$, the former exhibits a more intensive stress distribution, with higher values observed at various branches of visceral arteries. Moreover, the stress of Case 1 is significantly higher than that in the rest cases. This discrepancy is due to the strong material anisotropy for Case 1. The above numerical results reveal that, under physiologically realistic conditions, the strain and stress distribution may vary significantly due to the material models, and the impact of anisotropy is non-negligible when there is an interest in analyzing the deformation and stress of the vascular wall.
Figure 24: The von Mises stress of the arterial wall computed near the peak systole.
Figure 25: Magnitude images of 2D PCMRI at the moments of maximum (a) and minimum (b) cross-sectional areas.
Case Area variation range (cm2) Relative change in area
@@
PCMRI $1.85\sim 2.04$ 10.27%
Table 2: A comparison of the area variations between PCMRI measurements and FSI simulations at the slice indicated in Figure 18.
Figure 25 presents the 2D PCMRI images at the slice indicated in Figure 18. The regions of the lumen are delineated by blue closed curves, with the numbers indicating the areas of the enclosed regions. The area is calculated by counting the number of pixels. We further compare the area variations at this cross-section between the PCMRI measurement and FSI simulations, with the results listed in Table 2. The area variation range measured by PCMRI is in close agreement with predictions from the GOH model of $\kappa=0.226$. It implies that incorporating fiber dispersion within the GOH model offers a more accurate prediction of the behavior of the abdominal arterial wall.
@@ 4 Conclusion
In this work, we present a systematic methodology for modeling anisotropic arterial walls. Building upon existing vascular lumen modeling techniques, the arterial wall models are generated by manipulating segmentation data and employing CAD techniques. We introduce a morphology-based method to construct local basis vectors, which relies on the centerline information. In thick-walled geometries, the local basis vectors exhibit transmural variations, while in multi-branched vessels and near bifurcation regions, identifying the correct centerline poses challenging. Through a careful design, we proposed a complete pipeline for robustly generating local basis vectors for complex arterial geometries. Compared with methods that utilize elastic stress field to define local basis vectors, our approach has been demonstrated to deliver local basis vectors with improved quality. We additionally introduced a method for computing vascular prestress as a necessary initialization step for vascular FSI analysis.
@@
We analyze the influence of the anisotropy on biomechanical factors through two numerical examples. We find that the material properties of the arterial wall do not significantly impact on the hemodynamic factors, such as flow field and WSS. However, wall deformation and stress critically relate to the wall’s constitutive models. In particular, the deformation of isotropic and anisotropic wall models under physiological loading can be drastically different. Also, anisotropic models typically exhibit a localized stress distribution, while the isotropic model tend to develop a smoother stress field. Furthermore, fiber dispersion is highly relevant for the mechanical behavior of the wall, and a GOH model accounting for fiber dispersion shows high-quality agreement with the PCMRI measurements. This underscores the importance of a physiologically detailed arterial wall model in patient-specific investigations.
@@
Several areas for future research warrant further investigation. First, the arterial wall has a multi-layered structure, with distinct material properties and fiber orientations in the media and adventitia layers. Future research will develop models that accurately represent this layered structure and explore the biomechanical behavior of such layered models. Second, refining the models to better simulate the mechanical behavior of fibers is essential. This includes considering out-of-plane dispersions and excluding fiber responses under compression. These refinements will contribute to a more accurate representation of the arterial wall’s mechanical response and assist in better understanding its biomechanics and mechanobiology properties.

### v2

@@ 3.3 Patient-specific abdominal aorta model
The images at the cross-sections (illustrated by the dashed lines in Figure 18 (a)) of the visceral arteries, supra-celiac (SC) aorta, and infrarenal (IR) aorta are obtained using a pulse and respiratory gated 2D phase contrast sequence. The obtained 2D phase-contrast magnetic resonance imaging (PC-MRI) data are processed using Segment [1] to extract the physiological flow rates at these cross-sections for 30 time frames within a cardiac cycle. Polynomial interpolation and phase correction are applied to these flow rates to obtain the flow waveforms at the inlet and outlets through a cardiac cycle. Due to limitations in spatial resolution, the 2D PC-MRI technique is insufficient to accurately identify small arteries, such as the left gastric arteries and lumbar arteries, their flow rates are excluded. To ensure mass conservation, the sum of the flow waveforms of the IR aorta and all visceral arteries is set as the flow waveform for the inlet of the SC aorta. These obtained flow waveforms are then combined with a parabolic flow profile to prescribe the velocity at the inlet and the outlets of the five visceral arteries (Figure 18 (b)).
Figure 19: The mesh of AA model.
@@
Regarding the outlet of the IR aorta, a three-element Windkessel model is coupled with the 3D FSI model as the outflow boundary condition. Utilizing the measured flow waveform at the IR arterial outlet, the RCR parameters are tuned to ensure that the peak and trough of the output pressure match the systolic pressure ($P_{\mathrm{sys}}=103$ mmHg) and diastolic pressure ($P_{\mathrm{dia}}=72$ mmHg) of the subject. The specific settings are as follows. The compliance of the Windkessel model is set to 0.000452 cm3/dyn; the proximal and distal resistances are set to 330.69 $\cdots$/cm3 and 7708.28 $\cdots$/cm3, respectively. In this study, the exterior wall of the artery is kept stress-free, meaning the supporting tissue is not considered. Meanwhile, on the inlet and outlet solid surfaces, the ring-shaped walls are fixed.
We employ the isotropic neo-Hookean model (2.4) and the anisotropic GOH model (2.5) for the arterial wall to perform a comparative study. The solid material properties utilized are adopted from [9] and are summarized in Table 1. Moreover, the effect of fiber dispersion is investigated by considering two different dispersion parameters (i.e., $\kappa=0.0$ and $0.226$ [38]). The blood density is set to 1.06 g/cm3, and the dynamic viscosity is taken to be 0.04 $\cdots$/cm2. The geometric model is discretized into $8.52\times 10^{6}$ linear tetrahedral elements, including $4.64\times 10^{6}$ solid and $3.89\times 10^{6}$ fluid elements. Three layers of BL elements with a growth ratio of 0.75 are constructed within the lumen. In the meantime, we strive to have at least six layers of elements across the vessel wall thickness (Figure 19). We also investigate the problem with a relatively coarse mesh, which delivers essentially identical results. To construct the orientation of fibers, the local basis vectors are generated on the vessel wall using the procedures developed in Section 2.2 (Figure 20).
@@
Figure 21: The velocity computed at peak systole.
FSI simulations are performed using the initialization procedure outlined in Section 2.4. Two cardiac cycles are simulated with a fixed time step size of $\Delta t=1.0\times 10^{-5}$ s, and the results from the last cycle are presented here. Figures 21 and 22 depict the flow fields and WSS obtained by the four different material models at peak systole. It is evident that the flow and WSS distributions derived from these models are quite close, with only minor discrepancies in the flow field near the outlet of the IR aorta.
Figure 22: The WSS computed at peak systole.
Figure 23: The maximum principal strain of the arterial wall computed at peak systole.
Figures 23 and 24 depict the maximum principal strain and von Mises stress at peak systole for the four material models, respectively. For the isotropic models, the principal strain is lower and distributed smoothly within the wall; for the anisotropic models, the principal strain is higher in the visceral arteries. The stress pattern exhibits similarity between the two neo-Hookean models and the GOH model with $\kappa=0.226$. Comparing the stress results for the GOH model with $\kappa=0.0$ and $\kappa=0.226$, the former exhibits a more localized stress distribution, with higher values observed at various branches of visceral arteries. Moreover, the stress of Case 1 is significantly higher than that in the rest cases. This discrepancy is due to the strong material anisotropy for Case 1.
Figure 24: The von Mises stress of the arterial wall computed at peak systole.
Figure 25: Magnitude images of 2D PC-MRI at the moments of maximum (a) and minimum (b) cross-sectional areas.
Case Area variation (cm2) Relative change in area
@@
PC-MRI $1.85\sim 2.04$ 10.27%
Table 2: A comparison of the area variations between PC-MRI measurements and FSI simulations at the slice indicated in Figure 18.
Figure 25 presents the 2D PC-MRI images at the slice indicated in Figure 18. The regions of the lumen are delineated by blue closed curves, with the numbers indicating the areas of the enclosed regions. The area is calculated by counting the number of pixels. We further compare the area variations at this cross-section between the PC-MRI measurement and FSI simulations, with the results listed in Table 2. The area variation measured by PC-MRI is in close agreement with predictions from the GOH model of $\kappa=0.226$. The rest models give an over-stiff behavior of the wall, making the relative area change significantly less than the measurement. This signifies that incorporating fiber dispersion within the GOH model offers a more accurate prediction of the behavior of the abdominal arterial wall.
@@ 4 Conclusion
In this study, we first presented a systematic modeling pipeline for fibrous arterial walls in vascular FSI analysis. We proposed modeling the wall based on segmentation data and CAD tools. It is constructed upon the mature lumen modeling strategy, rendering it readily integrable into the mainstream cardiovascular modeling pipeline [7, 76]. In comparison with the tissue model generated via the extrusion procedure [51, 60, 79], the proposed strategy addresses the potential self-intersection issue for multi-branched geometries in a robust and flexible manner. Moreover, we have developed a specialized mesh generation procedure that allows generating BL mesh for the fluid subdomain and ensures mesh continuity across the luminal interface. Leveraging the centerline information, we introduced a morphology-based method to construct local basis vectors with the goal of facilitating the accurate description of fiber orientation in the image-based vascular geometric model. Specialized procedures have been developed to effectively handle thick-wall geometries, multi-branched vessels, and bifurcation regions. Compared with methods that utilized elastic stress field to define local basis vectors, our approach demonstrates improved quality in delivering local basis vectors.
@@
We analyzed the biomechanical significance of anisotropy through two numerical examples. Our findings indicate that the wall deformation and stress of neo-Hookean and GOH wall models under physiological loading are drastically different. By adjusting the dispersion parameter and fiber orientation, we confirmed that the difference arises from anisotropy instead of the exponential energy form. In both examples, we also observed that anisotropic models tend to deliver a localized stress distribution, whereas the isotropic models develop smoother stress fields. We identified that the stress distribution in the GOH model correlates with $\mathcal{I}$, the term appearing in the exponential function of the free energy. Furthermore, a GOH model accounting for fiber dispersion shows high-quality agreement with PC-MRI measurements, underscoring the importance of physiologically detailed arterial wall models in patient-specific investigations. The results indicate that the impact of anisotropy is non-negligible when analyzing the biomechanical behavior of the wall. On the other hand, we also observed that hemodynamic factors, such as WSS and flow field, are less sensitive to the choice of the wall model. Therefore, it may be acceptable to adopt a simple isotropic model when the focus is solely on the blood flow.
@@
Several areas for future research warrant further investigation. First, the proposed pipeline can be conveniently generalized to specify layered structure within the tissue, allowing the description of distinct material properties and fiber orientations in the media and adventitia layers. Future research will focus on developing models that accurately represent this layered structure and exploring the biomechanical behavior of such layered models, particularly in the FSI setting. Second, refining the models to better simulate fibers is essential. This includes considering out-of-plane dispersions and excluding fiber responses under compression. These refinements will contribute to a more accurate representation of the arterial wall and assist in better understanding its biomechanics and mechanobiology properties.
@@ Acknowledgements
Ethical Statement
Ethical statement is not applicable for this article.